OpenStudy (anonymous):

Solve the problem. After receiving a discount of 15.5% on its bulk order of typewriter ribbons, John's Office Supply pays $3549. What was the price of the order before the discount?

OpenStudy (cornzyblack):

(15.5/100)x = 3549 so x = (3549)/(0.155) x = 22896.77

OpenStudy (anonymous):

i dont think so :((

OpenStudy (cornzyblack):

I tried my best

OpenStudy (anonymous):

thanks